Differences between revisions 48 and 49
Revision 48 as of 2025-03-14 11:39:57
Size: 28144
Comment:
Revision 49 as of 2025-03-15 11:43:32
Size: 28138
Comment:
Deletions are marked like this. Additions are marked like this.
Line 194: Line 194:
The observed bump at ~0.3 Hz in array 3 is not evident now and the noise spectra have improved. Box C (array 2) KIDs (particularly KC037 and 025) still being a bit noisy. Screenshots from the DAQ are shown below. The observed bump at ~0.3 Hz in array 3 is not evident now and the noise spectra have improved. Box C (array 2) KIDs (particularly KC037 and 025) still being noisy. Screenshots from the DAQ are shown below.

Back to the NIKA2 Upgrade 2025 main page

Daily Reports

March 4th, Tuesday

Samuel, Alessandro, Martino and Gregory arrived at the telescope with many others from the IRAM Granada staff.

- First we take the decision that it is useless to check the laser alignment since it was done by Dave a few months ago and nothing was touched since then. We'll use Dave's measurement as reference.
- Today we dismounted the cryostat, with the help of Dave, removing the nose, moving the cryostat to install the rails allowing to slide all the screens to the wall, where was the nose, and access the 100mK structure where the clod optics stands (100mK lenses, dichroic, polariser, arrays). Pictures of all the steps steps are posted below.
At the end of the morning we reached the 4K screen.
On the afternoon we slid away the remaining screens, dismounted the dichroic (#2), the polariser, the smooth array side lenses of the 1mm channels, and the arrays.
- A bit of machining in the mechanical workshop was necessary to enlarge rings allowing to install the AR structured lenses. We installed these lenses and the new 1mm arrays. In parallel we changed the IR filters on the 300K screen, the 100K screen, and the 50K screen, that's the only ones we change given the type of filters provided by Cardiff (C7.5, C15, but no C30). The 3 new double faced filters we installed are all C15 types. We prepared the installation of the dichroic #1 new mount, with some machining also necessary to install the adaptation pieces on the 100 mK structure. We finished these step just before diner. We are a bit ahead of scheduled (originally we thought doing only the dismounting parts today).
- Tomorrow we install the dichroic with the new mount, then close everything back, which will take some time because of the super insulation to change at least on one screen. If everything goes well we might start pumping by the end of the day.

1) Opening NIKA2

250304_1_starting_dismounting_NIKA2 250304_02_NIKA2_back_before_displacement 250304_03_NIKA2_pedestal_before_displacement 250304_04_NIKA2_lifted_air_pressure 250304_05_NIKA2_pedestal_lifted 250304_06_NIKA2_displaced 250304_07_sliding_rails_insertion 250304_08_NIKA2_back_on_wall 250304_09_sliding_rails_in_place 250304_10_venting_nose 250304_11_removing_nose_head 250304_12_nose_100K_stage 250304_13_nose_100K_2nd-screen 250304_14_nose_50K_stage 250304_15_nose_measuring_50K_stage_height_right 250304_16_nose_measuring_50K_stage_height_left 250304_17_tie-rod_for_50K_stage_height_adjutment 250304_18_tie-rod_dismounted_1of2 250304_19_dismounting_50K_nose_stage_M7-M8 250304_20_50K_nose_stage_M7-M8_dismounted 250304_21_dismounting_nose_4K_stage 250304_22_cryostat_body_1K_stage_entrance 250304_23_sliding_300K_body_screen 250304_24_yellow_wheels_rails_inserted_in_300K_screen 250304_25_positioning_the_rolling_support_for_lower_T_screens 250304_26_rolling_support_bellow_100K_screen 250304_27_sliding_100K_screen 250304_28_inserting_yellow_wheels_rails_in_100K_screen 250304_29_sliding_50K_screen 250304_30_black_wheels_rails_to_insert_in_50K_screen 250304_31_4K_body_screen_not_yet_removed 250304_32_baffle_in_4K_nose_screen 250304_33_4K_and_1K_screens_removed 250304_34_removing_ice_from_the_100mK_stage 250304_35_100mK_stage 250304_36_1K_lens_through_the_100mK_magnetic_shield

2) Dismounting dichroic, arrays, lenses, filters to be changed, and installing the new ones

250304_37_removing_dichroic_2 250304_38_Array_3_removed 250304_39_Arrays_2_and_1_on_100mK_stage 250304_40_Arrays_3_and_1_removed_back_holder_side 250304_41_Arrays_3_and_1_removed_KIDs_side 250304_42_removing_polarizer 250304_43_removing_smooth_array_side_lenses 250304_44_installing_new_Array_3 250304_45_polarisation_mark_back_of_new_Array_3 250304_46_information_back_of_new_Array_3 250304_47_old_and_new_Array_1 250304_48_new_Array_1_KIDs_side 250304_49_installing_new_Array_1 250304_50_old_300K_IR_filter 250304_51_old_300K_IR_filter_and_entrance_window_AR_surf 250304_52_new_300K_IR_filter_installed 250304_53_old_100K_IR_filter 250304_54_new_100K_IR_filter_installed 250304_55_remounting_polarizer 250304_56_remounting_polarizer_anti_stray_light_black_box 250304_57_old_50K_IR_filter 250304_58_50K_nose_stage_M7_inside 250304_59_M7_inside_50K_nose_stage 250304_60_M8_inside_50K_nose_stage 250304_61_new_50K_IR_filter 250304_62_drilling_holes_for_adaptation_pieces_of_dichroic_1_new_mount 250304_63_drilling_holes_for_adaptation_pieces_of_dichroic_1_new_mount 250304_64_drilling_fixation_holes_in_dichroic_1_new_mount

March 5th, Wednesday

- The dichroic 1 must be mounted with a 0.5 N.m torque on the screws that allow holding in sandwich the dichroic membrane and its rings between the copper parts of its mount. Yesterday we found out we were missing the dynamometric wrench allowing to get this relatively low torque. Carlos called Granada to find the one he had on the lab down there, and this morning Antonio brought it up to the telescope. In the meantime we prepared the mount. A bit of additional machining was necessary in order to install the dichroic already mounted on NIKA2 as it is impossible to install the mount first and fix the dichroic afterward because of lack of space for the wrench. Once the dichroic installed we started closing the cryostat just before lunch.
- After lunch we continued the closing process, screen after screen, checking on the way that the cold mirrors block was exactly at the same height as the one measured during the dismounting process. Around 16:30 we closed the last parts of the 300K assembly, and started pumping. While pumping, look with Carlos and Victor how to add removable wheels to the PIMP mount to ease its displacement ; Victor built the wheels system. Leak check on NIKA2 around 18:30. No leaks found. We then put back the cryostat on its original working place.
- Pumping for a few hours. We might start the cooling already tonight, otherwise tomorrow morning.

3) Continuation of the installation of the new elements (dichroic)

250305_65_dichroic_1_new_mount_installed 250305_66_outer_adaptation_piece_for_d1_new_mount 250305_67_inner_adaptation_piece_for_d1_new_mount 250305_68_removing_d1_new_mount 250305_69_installing_d1_in_new_mount_with_Al_flattening_ring 250305_70_applying_50cNm_torque_to_screws_linking_clamping_ring_to_mount_with_d1_in_between 250305_71_mounting_dichroic_1_on_NIKA2 250305_72_dichroic_1_mounted_on_NIKA2

4) Closing NIKA2, start pumping and leak check

250305_73_sliding_back_magnetic_shield 250305_74_sliding_back_1K_screen 250305_75_screwing_back_1K_screen 250305_76_sliding_back_4K_screen 250305_77_sliding_back_50K_screen 250305_78_screwing_back_100K_screen 250305_79_1K_filter_ans_superinsulation_layers_in_cryostat_imput 250305_80_4K_noze_box_back_in_place 250305_81_mounting_back_the_noze_50K_stage 250305_82_adjusting_tie_rods_for_M7-M8_50K_stage_height 250305_83_checking_M7-M8_50K_stage_height 250305_84_checking_M7-M8_50K_stage_height 250305_85_reinstalling_noze_100K_2nd_screen 250305_86_reinstalling_noze_100K_2nd_screen 250305_87_noze_100K_1nd_screen_and_1st-3rd_of_300K_shell_reinstalled 250305_88_screwing_noze_3rd-3rd_300K_shell 250305_89_pumping_and_leak_check_on_closed_NIKA2 250305_90_NIKA2_about_to_be_put_back_on_its_working_position 250305_91_cryostat_front_positioning_pin_to_be_inserted_in_its_hole 250305_92_cryostat_back_positioning_pin_in_its_hole_and_removing_2nd_roll_ball 250305_93_NIKA2_cryostat_back_on_its_original_position 250305_94_HWP_and_NIKA2_back_on_position

March 6th, Thursday

- NIKA2 cryostat under vacuum. Cooling down started before 9:00.
- Laser alignment checked, removing M3 and using the elevation bearing laser. The laser spot is on the crosses of M4, M5, M6, within 1mm uncertainty, about 1mm above the centre of the cryostat window, exactly like Dave's measurement a few month ago.
Note: if during commissioning we find out that the focal surface is asymmetric, with worse beams concentrated on a local outer zone of the Arrays, we might play on the cryostat reference level value to improve the position of the arrays on the focal surface and get something more symmetrical.
- We have finished everything before lunch, ahead of schedule. There was no major issue during the upgrade process, so we didn't have to use the contingency time we had foreseen in case of problem. We decide to go back down to Granada, and get back homme 1 day earlier than planned.

5) Check laser alignment

250306_95_elevation_bearing_laser_M3_side 250306_96_laser_spot_on_M4_optical_axis_mark 250306_97_laser_through_M3_mount_seen_on_M5_optical_insert 250306_98_laser_spot_on_M4_M5_M6_and_NIKA2_window_cap 250306_99_laser_spot_on_NIKA2_window_cap 250306_100_laser_spot_on_NIKA2_window

Bye bye NIKA2. Next episode is checking detectors are alive and resonances can be found, then do the initial sweep. This will happen at middle of next week. And if everything is fine, on the following week we start the commissioning, see Nika2Run72 !

March 7th-10th, Friday-Monday

NIKA2 is cooling down.

March 11th, Tuesday

NIKA2 has reached the 1K zone, start condensation of the 3He-4He mixture.

March 12th, Wednesday

Cryostat okay, finale: 1mm H amplifier switched on at 07:45 (DJ). T_MC=0.169K, T_Still=1.23K, T_4K=8.5. So, ready for the sweep.

New sweep done (08:30-10:00), first approach: AM, DJ, AB. Lines transmission okay. Bad KIDs (a1, a2, a3): 4/953, 12/578, 12/1000.

Array 3 time lines/noise traces show some glitches/a correlated noise feature at ~0.3 Hz (see screenshots below): PTs adjustment? Arrays 1 and 2 look okay.

Similar screenshots after the sweep in Oct. 2022 can be seen at the bottom of this site.

A new attempt for refining resonances tables will be carried out tomorrow afternoon or Friday morning. Cryo run in force: #84.

- DAQ monitor: main control & lost KIDs:

run84_monitor_lostp_20250312 run84_monitor_mc_20250312

- Time lines & noise traces (arrays 1, 2 & 3):

timelines_a1_20250312 noisetr_a1_20250312

timelines_a2_20250312 noisetr_a2_20250312

timelines_a3_20250312 noisetr_a3_20250312

March 14th, Friday

Telescope stopped, wind speed < 5 m/s; cryostat: T_MC=0.159K, T_Still=1.19K, T_4K=8.2.

New sweep try done (09:00-11:00): AM, MC, DJ, AB. Bad KIDs (a1, a2, a3): 10/953, 12/578, 9/1000.

Setting attenuation values again for array 3. Box S KIDs ones were retouched after a couple of sweeps. No changes in resonances tables of array 1 and 2 wrt. the previous version. DAQ restarted, level 6: attente.

The observed bump at ~0.3 Hz in array 3 is not evident now and the noise spectra have improved. Box C (array 2) KIDs (particularly KC037 and 025) still being noisy. Screenshots from the DAQ are shown below.

- DAQ monitor: main control & lost KIDs:

run84_monitor_lostp_20250314 run84_monitor_mc_20250314

- Time lines & noise traces (arrays 1, 2 & 3):

timelines_a1_20250314 noisetr_a1_20250314

timelines_a2_20250314 noisetr_a2_20250314

timelines_a3_20250314 noisetr_a3_20250314

DailyReportsNika2Upgrade2025 (last edited 2025-03-15 11:43:32 by NikaBolometer)